OpenStudy (anonymous):

I need someone to explain this to me Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v.

TheSmartOne (thesmartone):

\(\bf a <b,c>~ =~ <ab,ac>\)

OpenStudy (anonymous):

I don't understand?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

\(\bf y=~<a,b>\) \(\bf z=~<c,d>\) \(\bf y-z =\) \(\sf <a,b> - <c,d>~= ~<a,b> -<c,d>~=~<a-c,~b-d>\)

OpenStudy (anonymous):

okay im even more confused

TheSmartOne (thesmartone):

\(\bf y=~<a,b>\) \(\bf z=~<c,d>\) \(\bf y-z =\) \(\sf <a,b> - <c,d>~=~<a-c,~b-d>\) Oops, fixed the redundant error...

TheSmartOne (thesmartone):

https://www.mathsisfun.com/algebra/vectors.html Review the topic :p

OpenStudy (anonymous):

could you possibly plug in the numbers for the Y & Z?

OpenStudy (anonymous):

Yes i need an explanation, not symbols

TheSmartOne (thesmartone):

\(\bf u= ~<-5,1>\) \(\bf 9u = 9<-5,1>~= <9\times (-5) , 9\times 1>\)

TheSmartOne (thesmartone):

That's unfortunate, because we can't solve this without the vector symbol :p

TheSmartOne (thesmartone):

so what is \(\sf 9\times (-5) =?\) and \(\sf 9\times1 =?\)

OpenStudy (anonymous):

-45 & 9

TheSmartOne (thesmartone):

correct, so now we have \(\bf 9u =~<-45,9>\)